OpenStudy (anonymous):

FreddieMac reports that the average rate on a 30-year fixed rate mortgage is 3.92% as of January 2012. This is down from 4.76% in January 2011 and 5.03% in January 2010. If you have a $225,000, 5%, 30-year mortgage, how much interest will you save if you refinance your loan at 3.5% for 15 years? The amortization table has 5% for 30 years at 5.37 and 3.5% for 15 years at 7.15. I tried 144180.00, but that is the wrong answer. Please help!!

OpenStudy (anonymous):

The answer is $145395.00. 225000.00\1000 = 225.00 x 5.37 = 1208.25 1208.25 x 360 = 434970.00 - 225000.00 = 209970.00 225000.00/1000 = 225.00 x 7.15 = 1608.75 1608.75 x 180 = 289575.00 - 225000.00 = 64575.00 209970.00 - 64575.00 = 145395.00